# PaperDebugger: A Plugin-Based Multi-Agent System for In-Editor Academic Writing, Review, and Editing

### v1

@@ Abstract.
In-editor writing assistance, LLM agents, Multi-agent orchestration, Overleaf integration, Academic writing tools
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing Interactive systems and toolsccs: Computing methodologies Natural language processingccs: Information systems Information retrieval
@@ 1. Introduction
Large language models (LLMs) are increasingly used in assisting academic writing workflows, from brainstorming and outlining to line-level editing and reviewer-response drafting. Recent systems for human–AI co-writing and assisted feedback demonstrate that LLM-based suggestions can improve fluency and reduce mechanical writing effort at scale (lee2024designspace; liebling2025towards). Research on writing-support tools further shows that structured interventions can meaningfully improve writing efficiency and user experience (sarrafzadeh2020stageaware; mysore2024pearl; ingley2023leveraging).
Despite these developments, the majority of existing tools still operate outside the primary editing environment, requiring copy-and-paste workflows and fragmenting interaction history (wen2024overleafcopilot; lee2024designspace). This external workflow introduces context switching, breaks writing flow, and makes revision history difficult to preserve. Additionally, external tools provide limited revision provenance; feedback, applied changes, and reasoning disappear once the interaction window closes. Tools such as Writefull provide in-editor language suggestions but remain largely surface-level, offering limited transparency into applied changes (writefull).
To address these challenges, we present PaperDebugger, an in-editor LLM agent system that integrates directly into Overleaf, a widely used academic writing editor. Instead of treating the writing process and model interaction as separate workflows, the system enables critique, refinement, and revision to take place in context, inline, and tied to document structure. The system provides persistent interaction history, patch-based edits, and structure-aware feedback while preserving the continuity of writing. Technically, it is implemented as a Chrome extension that communicates with a Kubernetes-based backend using gRPC. The Model Context Control Protocol (MCP) acts as a lightweight extensibility layer, enabling modular functionality and future agent capabilities without altering the core architecture. PaperDebugger is fully implemented with over 24,000 lines of code.
@@ 2.1. Architecture Overview
Agent Layer. The agent layer supports two execution modes: prompt-template agents and workflow-based agents. Prompt-template agents are lightweight, single-shot LLM invocations defined by structured templates. They are designed for low-latency tasks such as grammar polishing. Workflow-based agents are declarative workflows that coordinate multiple LLM calls, tool executions, and validation steps. These workflows handle complex tasks like deep research, relevant paper retrieval, and full-document enhancing.
@@ 2.2. Agentic Design
Building on these two execution modes, PaperDebugger extends beyond a single-model rewriting tool through the XtraMCP architecture, a refined variant of MCP tailored for academic writing. XtraMCP exposes a suite of validated tools for literature search, affiliation lookup, and structured data extraction, and enforces our internal Pydantic-based schemas and internal consistency checks to minimize hallucinations. Concretely, three core MCP-powered components back the agents: (i) a low-latency embedding + LLM re-ranking pipeline that provides high-quality semantic retrieval and real-time literature lookup; (ii) a multi-step AI review pipeline, inspired by conference reviewing workflows like AAAI, that guides the Reviewer agent through targeted, segment-level critique; and (iii) XtraGPT (chen2025xtragptcontextawarecontrollableacademic). XtraGPT is a model suite tuned for academic writing, ensuring that suggested revisions are context-aware, properly scoped, and phrased in appropriate scholarly style, more details can be found in the paper (chen2025xtragptcontextawarecontrollableacademic).
@@ 3. Usage Analytics
We analyze anonymized telemetry from both the Chrome extension and backend between May – November 2025 to understand how PaperDebugger is used in real writing environments. In the following, we present the key statistics, and more usage analysis is in our project repository.
@@
Table 1 shows early usage signals. With 112 extension installs, 78 registered users, and 23 monthly active users, roughly 30% of all users remain active month-to-month. Users created 158 projects and 797 writing threads, indicating sustained multi-session usage rather than one-off experimentation. User reviews reflect strong satisfaction with the integrated workflow (“convenience”, “seamless”). At the same time, feedback highlights natural research value and limitations, such as domain-specific tone (“suggestions feel CS-like”) and performance drops on long documents.
@@
Chrome extension installs 112
Registered users 78
Active users (30-day) 23
Projects created (all-time) 158
Threads created (all-time) 797
Chrome store user rating 4.9 / 5
@@
Diff viewed 1073
Copy suggestion 375
Insert patch 359
@@ 4.1. In-Editor Editing and Patch
A common task in AI-assisted academic writing is to polish or refine text using AI agents. For example, when revising a conference submission, an author may encounter an unclear section title and highlight it directly within Overleaf to request a structured rewrite. As shown in Figure 3, the selected text is forwarded to the PaperDebugger panel, where the author initiates a critique or refinement request.
Once invoked, PaperDebugger launches a coordinated agent pipeline comprising a critique agent, an enhancer agent, and a patch generator. The system returns the results as before–after diffs that can be inspected and applied directly within the editor.
All candidate patches are displayed as inline previews. The author examines the rationale behind each suggestion, selects the preferred revision, and applies it with a single click. This workflow replaces the traditional copy-and-paste interaction pattern common to external LLM tools, instead providing a seamless, context-aware editing loop inside Overleaf.
Overall, this case demonstrates how PaperDebugger elevates a simple editing action into a transparent and agentic patch workflow that enhances clarity and structure while preserving an efficient, in-editor writing experience.
Figure 3. In-editor editing workflow in PaperDebugger. (1) Select a span of LaTeX in Overleaf. (2) Add the selection to the PaperDebugger panel. (3) Specify the critique request. (4) Trigger the agentic pipeline. (5) Review and apply the returned before–after patches.

### v2

@@ Abstract.
In-editor writing assistance; LLM agents; Multi-agent orchestration; Overleaf integration; Academic writing tools
journalyear: 2026copyright: ccconference: Companion Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Companion Proceedings of the ACM Web Conference 2026 (WWW Companion ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774905.3793122isbn: 979-8-4007-2308-7/2026/04ccs: Human-centered computing Interactive systems and toolsccs: Computing methodologies Natural language processingccs: Information systems Information retrieval
@@ 1. Introduction
Large language models (LLMs) are increasingly used in assisting academic writing workflows, from brainstorming and outlining to line-level editing and reviewer-response drafting. Recent systems for human–AI co-writing and assisted feedback demonstrate that LLM-based suggestions can improve fluency and reduce mechanical writing effort at scale (Lee and et al., 2024; Liebling and et al., 2025). Research on writing-support tools further shows that structured interventions can meaningfully improve writing efficiency and user experience (Sarrafzadeh and et al., 2020; Mysore and et al., 2024; Ingley and Pack, 2023).
Despite these developments, the majority of existing tools still operate outside the primary editing environment, requiring copy-and-paste workflows and fragmenting interaction history (Wen and et al., 2024; Lee and et al., 2024). This external workflow introduces context switching, breaks writing flow, and makes revision history difficult to preserve. Additionally, external tools provide limited revision provenance; feedback, applied changes, and reasoning disappear once the interaction window closes. Tools such as Writefull provide in-editor language suggestions but remain largely surface-level, offering limited transparency into applied changes (Writefull Team, 2024).
To address these challenges, we present PaperDebugger, an in-editor LLM agent system that integrates directly into Overleaf, a widely used academic writing editor. Instead of treating the writing process and model interaction as separate workflows, the system enables critique, refinement, and revision to take place in context, inline, and tied to document structure. The system provides persistent interaction history, patch-based edits, and structure-aware feedback while preserving the continuity of writing. Technically, it is implemented as a Chrome extension that communicates with a Kubernetes-based backend using gRPC. The Model Context Protocol (MCP) acts as a lightweight extensibility layer, enabling modular functionality and future agent capabilities without altering the core architecture. PaperDebugger is fully implemented with over 24,000 lines of code.
@@ 2.1. Architecture Overview
Architecture diagram showing five layers: presentation, backend, agent, protocol, and infrastructure.
@@
Agent Layer. The agent layer supports two execution modes: prompt-template agents and workflow-based agents. Prompt-template agents are lightweight, single-shot LLM invocations defined by structured templates. They are designed for low-latency tasks such as grammar polishing. Workflow-based agents are declarative workflows that coordinate multiple LLM calls, tool executions, and validation steps. These workflows handle complex tasks like deep research, relevant paper retrieval, and full-document enhancement.
@@ 2.2. Agentic Design
Building on these two execution modes, PaperDebugger extends beyond a single-model rewriting tool through the XtraMCP architecture, a refined variant of MCP tailored for academic writing. XtraMCP exposes a suite of validated tools for literature search, affiliation lookup, and structured data extraction, and enforces our internal Pydantic-based schemas and internal consistency checks to minimize hallucinations. Concretely, three core MCP-powered components back the agents: (i) a low-latency embedding + LLM re-ranking pipeline that provides high-quality semantic retrieval and real-time literature lookup; (ii) a multi-step AI review pipeline, inspired by conference reviewing workflows like AAAI, that guides the Reviewer agent through targeted, segment-level critique; and (iii) XtraGPT (Chen et al., 2025). XtraGPT is a model suite tuned for academic writing, ensuring that suggested revisions are context-aware, properly scoped, and phrased in appropriate scholarly style, more details can be found in the paper (Chen et al., 2025).
@@
PaperDebugger end-to-end workflow: The extension captures user actions and sends them to the PaperDebugger server, which coordinates built-in agents or specialized agents on the XtraMCP server.
@@ 3. Usage Analytics
We analyze anonymized telemetry from both the Chrome extension and backend between May 2025 – January 2026 to understand how PaperDebugger is used in real writing environments. In the following, we present the key statistics, and more usage analysis is in our project repository.
@@
Table 1 shows early usage signals. With 4116 extension installs, 2761 registered users, and 732 monthly active users. Users created 3106 projects and 7447 writing threads, indicating sustained multi-session usage rather than one-off experimentation. User reviews reflect strong satisfaction with the integrated workflow (“convenience”, “seamless”). At the same time, feedback highlights natural research value and limitations, such as domain-specific tone (“suggestions feel CS-like”) and performance drops on long documents.
@@
Chrome extension installs 4116
Registered users 2761
Active users (30-day) 732
Projects created (all-time) 3106
Threads created (all-time) 7447
Chrome store user rating 4.82 / 5
@@
Diff viewed 4593
Copy suggestion 3214
Insert patch 2697
@@ 4.1. In-Editor Editing and Patch
A common task in AI-assisted academic writing is polishing or refining text with AI agents. For example, when revising a conference submission, an author may encounter an unclear section title and highlight it directly within Overleaf to request a structured rewrite. As shown in Figure 3, the selected text is forwarded to the PaperDebugger panel, where the author initiates a critique or refinement request.
Once invoked, PaperDebugger runs a coordinated pipeline with critique agent, enhancer, and patch generation agents, returning before-after diffs that can be inspected and applied directly within the editor.
Candidate patches appear as inline previews, allowing the author to review the rationale behind each suggestion, selects the preferred revision, and applies it with a single click. This workflow replaces traditional copy-paste interactions with a seamless, context-aware editing loop inside Overleaf.
Overall, this case demonstrates how PaperDebugger transforms a simple editing action into a transparent and agentic patch workflow that enhances clarity and structure while preserving an efficient, in-editor writing experience.
Figure 3. In-editor editing workflow in PaperDebugger. (1) Select a span of LaTeX in Overleaf. (2) Add the selection to the PaperDebugger panel. (3) Specify the critique request. (4) Trigger the agentic pipeline. (5) Review and apply the returned before-after patches.
In-editor editing workflow in PaperDebugger. (1) Select a span of LaTeX in Overleaf. (2) Add the selection to the PaperDebugger panel. (3) Specify the critique request. (4) Trigger the agentic pipeline. (5) Review and apply the returned before–after patches.
@@ 4.2. Deep Research and Comparative Analysis
Example of an end-to-end research-use scenario supported by PaperDebugger. The system integrates XtraMCP (a) deep research, (b) related-paper retrieval, and (c) section enhancer to help authors understand, compare, and refine academic content within the editor.
@@ 5. Conclusion
Acknowledgements.
This research is supported by the National Research Foundation, Singapore and Infocomm Media Development Authority under its Trust Tech Funding Initiative. Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not reflect the views of National Research Foundation, Singapore and Infocomm Media Development Authority.